OpenStudy (anonymous):

5x2 – 45 = 0 solve by using square roots

OpenStudy (goformit100):

its easy to solve it

OpenStudy (e.cociuba):

isn't this supposed to look like this 5x^2-45=0??

OpenStudy (anonymous):

yes it is

OpenStudy (e.cociuba):

Solution for 5x^2-45=0 equation: Simplifying 5x2 + -45 = 0 Reorder the terms: -45 + 5x2 = 0 Solving -45 + 5x2 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '45' to each side of the equation. -45 + 45 + 5x2 = 0 + 45 Combine like terms: -45 + 45 = 0 0 + 5x2 = 0 + 45 5x2 = 0 + 45 Combine like terms: 0 + 45 = 45 5x2 = 45 Divide each side by '5'. x2 = 9 Simplifying x2 = 9 Take the square root of each side: x = {-3, 3}